OpenStudy (abbycross167):

Can someone please help me to answer this math question? Please? I will fan and give a medal... Please help me to simplify this Rational Expression - 8 3 ------ - ----- x^2-4 x-2

Vocaloid (vocaloid):

hint: start by factoring x^2 - 4

OpenStudy (abbycross167):

it would be (2 + x)(-2 + x) right?

OpenStudy (mertsj):

Yes. We would write that (x-2)(x+2)

OpenStudy (mertsj):

To add or subtract fractions, the denominators have to be the same. So change the second fraction so that it has the same denominator.

Vocaloid (vocaloid):

good, then we convert both fractions to the same denominator any ideas now?

OpenStudy (abbycross167):

So change 8 3 ------ - ----- x^2-4 x-2 into 8 3 ------ - ----- x-2 x-2

Vocaloid (vocaloid):

not quite, we can't do it like that

Vocaloid (vocaloid):

remember, in order to keep a fraction the same, we have to change the numerator and the denominator by the same operations

Vocaloid (vocaloid):

to go from x - 2 to x^2 - 4, we have to multiply by (x+2)

OpenStudy (mertsj):

\[\frac{8}{(x-2)(x+2)} - \frac{3}{x-2}\] \[\frac{8}{(x-2)(x+2)}-\frac{3}{(x-2)}\times \frac{x+2}{x+2}\]

Vocaloid (vocaloid):

so... 3 --- x - 2 becomes 3(x-2)(x+2) ------ x^2 - 4

OpenStudy (mertsj):

Can you take it from there?

OpenStudy (abbycross167):

what do I do after that? @Mertsj

OpenStudy (mertsj):

Do you see that both denominators are now the same?

OpenStudy (abbycross167):

yes sir

OpenStudy (mertsj):

So leave the denominators alone. Use the distributive property on the numerator of the second fraction.

OpenStudy (mertsj):

What did you get?

OpenStudy (abbycross167):

3 x^2-6 x ? @Mertsj

OpenStudy (mertsj):

3(x+2)= ?????

OpenStudy (mertsj):